OpenStudy (anonymous):

Essay: Show all work. A school designer wants to create a whiteboard with the optimal dimensions to enhance learning. It is determined that if one side of the whiteboard is x + 2y inches, the other side should be inches. Write an algebraic expression for the area of such a whiteboard, simplify it, and include correct units with your solution. So, I just want to make sure I have done this correctly..... (x+2y)(x^2+6xy-5y^2)= A= side 1 * side 2 Side 1= x+2y Side 2= x^2+6xy-5y^2 A=(x+2y)(x^2+6xy-5y^2) A=x^3+6x^2y-10y^3

OpenStudy (anonymous):

the other side should be inches.

OpenStudy (anonymous):

im thinking the expression didnt show due to copy paste if you look at the work, its shown there

OpenStudy (anonymous):

side 2?

OpenStudy (anonymous):

\[x^3+8 x^2 y+7 x y^2-10 y^3\] is right

OpenStudy (anonymous):

\[(x+2y)(x^2+6xy-5y^2)= x(x^2+6xy-5y^2)+2y(x^2+6xy-5y^2)\]

OpenStudy (anonymous):

by which i mean \[(x+2y)(x^2+6xy-5y^2)=x^3+8 x^2 y+7 x y^2-10 y^3\]

OpenStudy (anonymous):

in other words, everything but your final solution is correct

OpenStudy (anonymous):

ok, I'm really blonde here lol. where did the 8x2y come from?

OpenStudy (anonymous):

\[x(x^2+6xy−5y^2)+2y(x^2+6xy−5y^2)\] \[x^3+6x^2y−5xy^2+2x^2y+12xy^2−10y^3\] \[x^3+(6x^2y+2x^2y)−5xy^2+12xy^2−10y^3\]

OpenStudy (anonymous):

oh, ok I get it... thank you so much for breaking that down for me!!

OpenStudy (anonymous):

its just distribution and then simplify

OpenStudy (anonymous):

thanks again:)